OpenStudy (anonymous):

Let y=[(8x+2)^4][(x^3)−8)^7] Use logarithmic differentiation to determine the derivative.

OpenStudy (anonymous):

\[\ln\left( (8x+2)^4(x^3-8)^7\right)=4\ln(8x+2)+7\ln(x^3-8)\] now the derivative is easy

OpenStudy (anonymous):

 can you take it from here?